OpenStudy (anonymous):

Can Someone Help Me Please simplify (3/7)^3

OpenStudy (anonymous):

just multiply it out... (3/7)^3 = (3/7)*(3/7)*(3/7) = ???

OpenStudy (anonymous):

you sure i did and got 0.07871720116618076

OpenStudy (anonymous):

just leave it in fraction form.... no need for calculator answers....

OpenStudy (anonymous):

what would it be then

OpenStudy (anonymous):

would it be 27/343

OpenStudy (phi):

yes

OpenStudy (anonymous):

okay thank you